# MSI-NeRF: Linking Omni-Depth with View Synthesis through Multi-Sphere Image aided Generalizable Neural Radiance Field

### v1

@@ Abstract
Panoramic observation using fisheye cameras is significant in robot perception, reconstruction, and remote operation. However, panoramic images synthesized by traditional methods lack depth information and can only provide three degrees-of-freedom (3DoF) rotation rendering in virtual reality applications. To fully preserve and exploit the parallax information within the original fisheye cameras, we introduce MSI-NeRF, which combines deep learning omnidirectional depth estimation and novel view rendering. We first construct a multi-sphere image as a cost volume through feature extraction and warping of the input images. It is then processed by geometry and appearance decoders, respectively. Unlike methods that regress depth maps directly, we further build an implicit radiance field using spatial points and interpolated 3D feature vectors as input. In this way, we can simultaneously realize omnidirectional depth estimation and 6DoF view synthesis. Our method is trained in a semi-self-supervised manner. It does not require target view images and only uses depth data for supervision. Our network has the generalization ability to reconstruct unknown scenes efficiently using only four images. Experimental results show that our method outperforms existing methods in depth estimation and novel view synthesis tasks.
I Introduction
Omnidirectional imagery is an important tool for robotics, computer vision, and virtual reality (VR). Compared to normal pin-hole images with limited field-of-view (FoV), omnidirectional images include all viewing angles both horizontally and vertically. It creates a compact way to store information about an entire scene in just one camera shot. Traditionally, omnidirectional images are synthesized using multiple fisheye cameras or camera arrays [1, 2]. Seamless omnidirectional images can be generated by matching and stitching overlapping pixels between adjacent images. They are then saved in equirectangular projected images using spherical coordinate representation.
@@ I Introduction
[…]
With the development of computer vision and deep learning, some methods try to solve such problems. OmniMVS [3] extends the traditional multi-view stereo (MVS) method to panoramic views. It does not directly stitch images together but uses deep learning methods to match features between different cameras and returns the panoramic depth map. However, it does not consider occluded regions and is insufficient to support 6DoF rendering. OmniNeRF [4] proposes a method for 6DoF rendering panoramically. However, since it requires processed omnidirectional images and additional depth as input, it can not be implemented end-to-end. The need for per-scene optimization also limits its application in fields with high real-time requirements. MVSNeRF [5] achieves free view rendering with generalizable network. Nevertheless, it only works for object-level or front-facing scene reconstruction. No research still achieves the reconstruction of panoramic scenes using a single shot.
This paper aims to generate an omnidirectional representation that preserves the 3D information in the input multi-view fisheye images. We can obtain omnidirectional depth information from such representation and perform 6DoF perspective synthesis. As input, we collect images from four surrounding fisheye cameras. Then, we follow the general MVS pipeline to build a multi-sphere image (MSI) at predefined depth layers as a cost volume. A NeRF [6] is then constructed to represent the scene with an implicit function. It takes the interpolated features from the cost volume as an additional input to better utilize the learned geometry and texture. For network training, our method utilizes semi-self-supervision. The traditional training pipeline cannot meet the needs since mainstream panoramic datasets only contain depth ground truth and do not include target views for training. Our method utilizes input images to introduce color supervision into the network, enabling generalizable NeRF training using only depth ground truth. The outline of our method is shown in Fig. 1.
With our approach, we can efficiently render omnidirectional depth maps and novel view images. Vision-based panoramic depth has the advantages of lower cost, higher resolution, and larger FoV compared with LiDAR. Such depth data plays an important role in autonomous driving, robot perception, and scene reconstruction. For the field of virtual reality, being able to render 6DoF novel views eliminates the VR sickness caused by 3DoF rendering. It also benefits the development of human-machine remote collaboration and teleoperation.
@@
    We combine the depth estimation task with the novel view synthesis task. By building a semi-self-supervised framework, multi-task network training can be achieved only using the most common depth data supervision.
@@
    Our trained network can generalize across scenes and achieve an inference speed of approximately ten frames per second. In experiments, our method performs well on both tasks and achieves state-of-the-art results.
II Related Works
II-A Omnidirectional Depth Estimation
Depth estimation using multi-view images has long been a concern in computer vision and robotics. By stereo-matching colors or features from different viewing angles, depth information can be obtained through the MVS methods [7, 8].
Unlike the general MVS algorithms, which can only generate depth maps of a specific target perspective, omnidirectional depth estimation aims to obtain panoramic depth using cameras surrounding all $360^{\circ}$. It usually requires multiple large FoV fisheye cameras [9, 3, 10, 11] or spherical camera array [12, 13] as input. SweepNet [9] modifies the Plane Sweep method [11] into spherical coordinates and uses Semi-Global Matching (SGM) [14] to obtain a depth map. Its following work, OmniMVS [3], also uses the quad fisheye array but improves the matching module to a 3D convolution network. Sphere-Sweeping-Stereo [10] proposes to use adaptive spherical matching and cost aggregation to accelerate inference while maintaining a fine-grained depth map quality.
@@ II-A Omnidirectional Depth Estimation
II-B Omnidirectional Novel View Synthesis
Novel view synthesis means to continuously interpolate new views given a limited set of source views. NeRF [6] and multi-plane image (MPI) [15] methods perform view synthesis on object level and forward facing scenes by constructing implicit or explicit geometric proxy. Although high quality novel view images can be generated, the bounded scene region still limits their practical usage.
In order to extend novel view synthesis to unbounded scenes, methods using full space mapping [16, 17] and omnidirectional input [4, 18] have been developed. Mip-NeRF 360 [16] enhances Mip-NeRF [19] with scene parameterization to realize $360^{\circ}$ scene representation. But it still needs densely captured views for high quality rendering. Methods taking omnidirectional images alleviates such problem. MatryODShka [18] takes omnidirectional stereo (ODS) images as input and uses neural network to blend them into MSI for novel view synthesis. OmniNeRF [4] first synthesizes multi-view images from color and depth panorama by point projection. It then trains a scene specific NeRF using the two-stage method, which lacks practicality.
@@ II-B Omnidirectional Novel View Synthesis
Figure 2: Structure of our method. Our method can be divided into three parts. First, by 2D feature extraction and warping, a multi-sphere image (MSI) representation can be built. Then, through geometric and appearance 3D decoder, explicit features ($\mathbf{f}_{geo}$, $\mathbf{f}_{appr}$) containing prior information can be obtained. They are then fed into NeRF implicit MLP along with point position $\mathbf{x}$, ray direction $\mathbf{d}$, and projected color $\mathbf{c}_{proj}$. The output occupancy and color are used for fisheye color image and depth image rendering. The learnable parameters in CNN and MLP are trained through semi-self-supervision.
II-C Generalizable NeRF
NeRF and its derived algorithms have shown great performance in 3D scene representation. However, most of them require scene-specific training. Whenever faced with a new scenario, these methods need seconds or even hours of training before they can generate high-quality results [20]. The nature of training from scratch also makes them dependent on dense source views to ensure that novel views can be correctly interpolated.
Leveraging the development of deep learning and stereo matching, NeRF with generalizability gradually emerged [21, 22, 5]. These generalizable NeRFs eliminate scene-specific training and can quickly synthesize novel views after training on large-scale datasets, even with only sparse view inputs. PixelNeRF [21] first uses a CNN encoder to extract image features from source views. It then builds an implicit function that takes the projected 2D features as additional input, following the idea of image-based rendering. IBRNet [22] proposes a similar way to achieve generalizability. It uses a ray transformer to add interaction along the ray during the render process. Differently, MVSNeRF [5] incorporates MVS methods into NeRF rendering. It first constructs a cost volume from the source views using plane sweeping and 3D convolution. Interpolated 3D features and the projected colors are then added to the NeRF input, transferring prior knowledge into the implicit field.
@@ II-C Generalizable NeRF
III Method
In this work, we present MSI-NeRF, a deep learning method that inputs four fisheye images and outputs a panoramic scene representation. It adopts the representation form of NeRF and uses priorly generated multi-sphere images as a geometric aid. Such representation enables 6DoF view synthesis while retaining depth information. During optimization, our method only requires depth ground truth. It can introduce color information into the network through input data to achieve semi-self-supervision. Our MSI-NeRF consists of three blocks: multi-sphere image construction (Sec. III-A), hybrid neural rendering (Sec. III-B), and multi-task semi-self-supervision (Sec. III-C). Its overall structure is shown in Fig. 2.
III-A Multi-Sphere Image Construction
@@
$\mathbf{V}_{n,i}(\theta,\phi)=F_{i}\left(\Pi_{i}\left(\overline{\mathbf{p}}(\theta,\phi)/d^{-1}_{n}\right)\right),$ (2)
where $\overline{\mathbf{p}}$ represents the unit ray for the spherical coordinate $<\theta,\phi>$ of the equirectangular projection, $\Pi_{i}$ is the projection function from rig coordinate to the $i^{th}$ camera, and $d^{-1}_{n}$ is the inverse depth of the $n^{th}$ layer. $\mathbf{V}$ alias the generated spherical volume representation, and $F$ is the image feature map.
@@
where $h_{geo}$ and $h_{app}$ represents the two decoders, and $\mathbf{V}_{geo}$ and $\mathbf{V}_{app}$ are the generated feature volumes. By aggregating the feature vectors from different views and conducting information interaction with neighboring spatial voxels, reasonable space occupancy, and color can be encoded in the resulting 3D feature grids $\mathbf{V}_{geo}$ and $\mathbf{V}_{app}$. They will be used as explicit aids for radiance field construction and neural rendering. By distinguishing between appearance and geometry volume representations, we can make our decoder network functions more concrete, each performing their own duties. It also allows interaction and clarity of network learning.
III-B Hybrid Neural Rendering
@@
$\mathbf{c}_{proj,i}$ $=I_{i}\left(\Pi_{i}\left(\mathbf{x}\right)\right),$
where $S$ represents the interpolation operation. In particular, we first apply positional encoding on the position and direction vectors. Then, we use an MLP taking embedded $\mathbf{x}$, $\mathbf{d}$, and $\mathbf{f}_{geo}$ as input, generating the occupancy and a feature vector. Finally, the feature vector is then processed by another MLP along with $\mathbf{f}_{appr}$ and $\mathbf{c}_{proj}$, generating the final color output.
Since we use inverse depth as the MSI layer arrangement criterion, for distant layers, the distance between two adjacent layers is also greater. In this way, using 3D interpolation will cause wrongly mixed prior features and affect convergence during training. As a result, we use the intersection of rays and spherical depth layers as our sample points. This way, only 2D interpolation on the 3D volume is needed, avoiding blending between depth layers with large gaps while maintaining balanced sample points. The intersection is calculated as follows:
@@
where $\mathbf{r}_{o}$ and $\mathbf{r}_{d}$ are the origin and unit direction of the ray $\mathbf{r}$, and $z_{n}$ is the depth of the intersecting point with the $n^{th}$ sphere along the ray. Note that the equation only applies to the case where rays are emitted inside the sphere. After points sampling, we follow the pipeline of UNISURF [23] to conduct volume rendering:
@@
$w\left(\mathbf{x}_{i}\right)=o\left(\mathbf{p}_{i}\right)T\left(\mathbf{p}_{i}\right),\quad T\left(\mathbf{p}_{i}\right)=\prod_{j=1}^{i-1}\left(1-o\left(\mathbf{p}_{j}\right)\right).$ (9)
@@
III-C Multi-Task Semi-Self-Supervision
In order to let our network generalize across different scenes, training on large-scale panoramic image datasets is needed. However, large-scale datasets containing source fisheye images as input and multi-view target images as ground truth are difficult to obtain. Such difficulties will lead to trouble in learning scene appearance and novel view synthesis. In this work, we utilize the widely used fisheye omnidirectional depth estimation datasets proposed by OmniMVS [3] for both geometry and appearance understanding of panoramic scenes. We revisit the training mechanism of generalizable NeRFs and find that the purpose of using multi-view target images for network training is: (a) to predict scene occupy information and (b) to predict the appearance information of the occluded part of the scene. These two points allow generalizable NeRFs to quickly synthesize novel views from a small number of target views, using the learned prior knowledge. Now that we already have the depth supervision to learn occupancy, we only need to find a way to inject occluded color supervision into training. We come up with the idea of directly using the input fisheye images as a multi-view supervision to provide texture information for the occluded parts.
To this end, we propose our multi-task semi-self-supervision strategy. For the original depth ground truth aligned with the rig coordinate, we use the rendered depth to compare it as depth supervision. We further render fisheye color images using the source camera’s intrinsic and extrinsic parameters and compare them with the captured images, forming an additional color supervision. Our loss function is represented as:

### v2

@@ Abstract
Panoramic observation using fisheye cameras is significant in virtual reality (VR) and robot perception. However, panoramic images synthesized by traditional methods lack depth information and can only provide three degrees-of-freedom (3DoF) rotation rendering in VR applications. To fully preserve and exploit the parallax information within the original fisheye cameras, we introduce MSI-NeRF, which combines deep learning omnidirectional depth estimation and novel view synthesis. We construct a multi-sphere image as a cost volume through feature extraction and warping of the input images. We further build an implicit radiance field using spatial points and interpolated 3D feature vectors as input, which can simultaneously realize omnidirectional depth estimation and 6DoF view synthesis. Leveraging the knowledge from depth estimation task, our method can learn scene appearance by source view supervision only. It does not require novel target views and can be trained conveniently on existing panorama depth estimation datasets. Our network has the generalization ability to reconstruct unknown scenes efficiently using only four images. Experimental results show that our method outperforms existing methods in both depth estimation and novel view synthesis tasks.
1 Introduction
Omnidirectional imagery is an important tool for computer vision, VR, and robotics. Compared to normal pin-hole images with limited field-of-view (FoV), omnidirectional images include all viewing angles both horizontally and vertically. It creates a compact way to store information about an entire scene in just one camera shot. Traditionally, omnidirectional images are synthesized using multiple fisheye cameras or camera arrays [5, 25]. Seamless omnidirectional images can be generated by matching and stitching overlapping pixels between adjacent images. They are then saved in equirectangular projected images using spherical coordinate representation.
Figure 1: Our method uses the images captured from four fisheye cameras arranged in a panoramic configuration as input. Aided by a multi-sphere image, a generalizable omnidirectional radiance field can be produced. From the radiance field, we can query occupancy and color information of any spatial position and ray direction. Then, leveraging the volume rendering formula, novel view synthesis, and depth estimation can be accomplished.
@@ 1 Introduction
With the development of computer vision and deep learning, methods have tried to solve such problems. OmniMVS [38] extends the traditional multi-view stereo (MVS) method to panoramic views. It uses deep learning methods to match features between different cameras and returns the panoramic depth map. However, it does not consider occluded regions and is insufficient to support 6DoF rendering. OmniNeRF [20] proposes a method for 6DoF rendering panoramically. However, since it requires processed omnidirectional images and additional depth as input, it can not be implemented end-to-end. The need for per-scene optimization also limits its application. MVSNeRF [9] achieves free view rendering with generalizable network. Nevertheless, it only works for object-level or front-facing scene reconstruction. No research has achieved the reconstruction of panoramic scenes using a single shot.
This paper aims to generate an omnidirectional representation that preserves the 3D information in the input multi-view fisheye images. We can obtain omnidirectional depth information from such representation and perform 6DoF perspective synthesis. As input, we collect images from four surrounding fisheye cameras, and follow the general MVS pipeline to build a multi-sphere image (MSI) at predefined depth layers as a cost volume. A NeRF [29] is then constructed to represent the scene implicitly. It takes the interpolated features from the cost volume as an additional input to better utilize the learned geometry and texture. Traditional novel view synthesis task usually needs datasets containing both source and target view images for training. But mainstream panoramic datasets only contain depth data and rarely include target views. Our method utilizes input images to introduce color supervision into the network, enabling generalizable NeRF training using only depth ground truth. The outline of our method is shown in Fig. 1.
With our approach, we can efficiently render omnidirectional depth maps and novel view images. Vision-based panoramic depth has the advantages of lower cost, higher resolution, and larger FoV compared with LiDAR. Such depth data plays an important role in autonomous driving, scene reconstruction and robotics application. For the field of virtual reality, being able to render 6DoF novel views eliminates the VR sickness caused by 3DoF rendering. It also proposes a new way of spatial video acquisition and editing, and can perform video de-shaking for both rotation and translation.
@@
    We combine the depth estimation task with the novel view synthesis task. By leveraging the inductive bias from MSI representation, multi-task network training can be achieved only using the most common depth data supervision.
@@
    Our trained network can generalize across scenes by pre-training on synthetic datasets. In experiments, our method performs well on both tasks and achieves state-of-the-art results. It can be used in various VR and robotics applications to eliminate VR sickness, enable panoramic video editing, and 3D reconstruction.
2 Related Works
2.1 Omnidirectional Depth Estimation
Depth estimation using multi-view images has long been a concern in computer vision [39, 38, 40] and robotics [14, 15, 41, 27]. By stereo-matching colors or features from different viewing angles, depth information can be obtained through the MVS methods [7, 42].
Unlike the general MVS algorithms, which can only generate depth maps of a specific target perspective, omnidirectional depth estimation aims to obtain panoramic depth using cameras surrounding all $360^{\circ}$. It usually requires multiple large FoV fisheye cameras [39, 38, 28, 17, 40, 10] or spherical camera array [46, 6] as input. SweepNet [39] modifies the Plane Sweep method [17] into spherical coordinates and uses Semi-Global Matching (SGM) [19] to obtain a depth map. Its following work, OmniMVS [38], also uses the quad fisheye array but improves the matching module to a 3D convolution network. Sphere-Sweeping-Stereo [28] proposes to use adaptive spherical matching and cost aggregation to accelerate inference while maintaining a fine-grained depth map quality. OmniVidar [40] improves the traditional fisheye camera model and proposes an attention-based recurrent network structure for high-quality depth estimation.
@@ 2.1 Omnidirectional Depth Estimation
2.2 Omnidirectional Novel View Synthesis
Novel view synthesis means to continuously interpolate new views given a limited set of source views. NeRF [29] and multi-plane image (MPI) [45, 24] methods perform view synthesis on object level and forward facing scenes by constructing implicit or explicit geometric proxy. Although high quality novel view images can be generated, the bounded scene region still limits their practical usage.
In order to extend novel view synthesis to unbounded scenes, methods using full space mapping [3, 44, 36, 11, 4] and omnidirectional input [20, 1, 22, 16, 35, 18, 30] have been developed. Mip-NeRF 360 [3] enhances Mip-NeRF [2] with scene parameterization to realize $360^{\circ}$ scene representation. But it still needs densely captured views for high quality rendering. Methods taking omnidirectional images alleviates such problem. MatryODShka [1] takes omnidirectional stereo (ODS) images as input and uses neural network to blend them into MSI for novel view synthesis. OmniNeRF [20] first synthesizes multi-view images from color and depth panorama by point projection. It then trains a scene specific NeRF using the two-stage method, which lacks practicality.
@@ 2.2 Omnidirectional Novel View Synthesis
Figure 2: Structure of our method. Our method can be divided into three parts. First, by 2D feature extraction and warping, a multi-sphere image (MSI) representation can be built. Then, through geometric and appearance 3D decoders, explicit features ($\mathbf{f}_{geo}$, $\mathbf{f}_{appr}$) containing prior information can be obtained. They are then fed into NeRF implicit MLP along with point position $\mathbf{x}$, ray direction $\mathbf{d}$, and projected color $\mathbf{c}$. The output occupancy and color are used for fisheye color image and depth image rendering, forming the final supervision loss.
2.3 Generalizable NeRF
NeRF and its derived algorithms have shown great performance in 3D scene representation. However, most of them require scene-specific training. Whenever faced with a new scenario, these methods need seconds or even hours of training before they can generate high-quality results [29, 31, 8]. The nature of training from scratch also makes them dependent on dense source views to ensure that novel views can be correctly interpolated.
Leveraging the development of deep learning and stereo matching, NeRF with generalizability gradually emerged [43, 37, 9, 34, 13, 12, 23, 26]. These generalizable NeRFs eliminate scene-specific training and can quickly synthesize novel views after training on large-scale datasets, even with only sparse view inputs. PixelNeRF [43] first uses a CNN encoder to extract image features from source views. It then builds an implicit function that takes the projected 2D features as additional input, following the idea of image-based rendering. IBRNet [37] proposes a similar way to achieve generalizability. It uses a ray transformer to add interaction along the ray during the render process. Differently, MVSNeRF [9] incorporates MVS methods into NeRF rendering. It first constructs a cost volume from the source views using plane sweeping and 3D convolution. Interpolated 3D features and the projected colors are then added to the NeRF input, transferring prior knowledge into the implicit field.
@@ 2.3 Generalizable NeRF
3 Method
In this work, we present MSI-NeRF, a deep learning method that inputs four fisheye images and outputs a panoramic scene representation. It adopts the representation form of NeRF and uses priorly generated multi-sphere images as a geometric aid. Such representation enables 6DoF view synthesis while retaining depth information. During optimization, our method only requires depth ground truth. It can introduce color information into the network through input data to achieve multi-task supervision. Our MSI-NeRF consists of three blocks: multi-sphere image construction (Sec. 3.1), hybrid neural rendering (Sec. 3.2), and multi-task supervision (Sec. 3.3). Its overall structure is shown in Fig. 2.
3.1 Multi-Sphere Image Construction
@@
$\mathbf{V}_{n,k}(\theta,\phi)=F_{k}\left(\Pi_{k}\left(\overline{\mathbf{p}}(\theta,\phi)/d^{-1}_{n}\right)\right),$ (2)
where $\overline{\mathbf{p}}$ represents the unit ray for the spherical coordinate, $\theta,\phi$ are the corresponding azimuth and elevation, $\Pi_{k}$ is the projection function from rig coordinate to the $k^{th}$ camera, and $d^{-1}_{n}$ is the inverse depth of the $n^{th}$ layer. $\mathbf{V}$ alias the generated spherical volume representation, and $F$ is the image feature map.
@@
where $h_{geo}$ and $h_{app}$ represents the two decoders, and $\mathbf{V}_{geo}$ and $\mathbf{V}_{app}$ are the generated feature volumes. They encode the spatial occupancy and color information respectively, and can be used as explicit aids for radiance field construction and neural rendering. By distinguishing between appearance and geometry volume representations, we can make our decoder network functions more concrete, each performing their own duties. It also allows interaction and clarity of network learning.
3.2 Hybrid Neural Rendering
@@
$\mathbf{c}_{proj,k}$ $=I_{k}\left(\Pi_{k}\left(\mathbf{x}\right)\right),$
where $S$ represents the interpolation operation, and $I_{k}$ is the $k^{th}$ color image. In particular, we first apply positional encoding on the position and direction vectors. Then, we use an MLP taking embedded $\mathbf{x}$, $\mathbf{d}$, and $\mathbf{f}_{geo}$ as input, generating the occupancy and a feature vector. Finally, the feature vector is then processed by another MLP along with $\mathbf{f}_{appr}$ and $\mathbf{c}_{proj}$, generating the final color output.
Since we use inverse depth as the MSI layer arrangement criterion, for distant layers, the distance between two adjacent layers is also greater. In this way, using 3D interpolation will cause wrongly mixed prior features and affect convergence during training. As a result, we use the intersection of rays and spherical depth layers as our sample points. This way, only 2D interpolation on the 3D volume is needed, avoiding blending between depth layers with large gaps while maintaining balanced sample points. The intersection is calculated by solving a quadratic equation:
@@
where $a$, $b$, and $c$ are the corresponding coefficients, and $\mathbf{r}_{o}$ and $\mathbf{r}_{d}$ are the origin and unit direction of the ray $\mathbf{r}$. The solution that conforms to the perspective relationship can be obtained by the root-finding formula:
@@
where $z_{n}$ is the depth of the intersecting point with the $n^{th}$ sphere along the ray. After points sampling, we follow the pipeline of UNISURF [32] to conduct volume rendering:
@@
$w\left(\mathbf{p}_{i}\right)=o\left(\mathbf{p}_{i}\right)T\left(\mathbf{p}_{i}\right),\quad T\left(\mathbf{p}_{i}\right)=\prod_{j=1}^{i-1}\left(1-o\left(\mathbf{p}_{j}\right)\right).$ (9)
@@
3.3 Multi-Task Supervision
In order to let our network generalize across different scenes, training on large-scale panoramic image datasets is needed. However, large-scale datasets containing both source fisheye images and multi-view target images are difficult to obtain. Such difficulties will lead to trouble in learning scene appearance and novel view synthesis. In this work, we utilize the widely used fisheye omnidirectional depth estimation datasets proposed by OmniMVS [38] for both geometry and appearance understanding of panoramic scenes. We revisit the training mechanism of generalizable NeRFs and find that the purpose of using multi-view target images for network training is: (a) to predict scene occupy information and (b) to predict the appearance information of the occluded part of the scene. These two points allow generalizable NeRFs to quickly synthesize novel views from a small number of target views, using the learned prior knowledge. Now that we already have the depth supervision to learn occupancy, we only need to find a way to inject occluded color supervision into training. We come up with the idea of directly using the input fisheye images as a multi-view supervision to provide texture information for the occluded parts.
To this end, we propose our multi-task supervision strategy. For the original depth ground truth aligned with the rig coordinate, we use the rendered depth to compare it as depth supervision. We further render fisheye color images using the source camera’s intrinsic and extrinsic and compare them with the captured images, forming an additional color supervision. Our loss function is represented as: